OpenStudy (anonymous):

What is the difference between Initial and Boundary value problems? What is the interval of validity for the following Initial Value Problem. 6(dy/dx) -2y=xy^4; y(0)=-2

OpenStudy (amistre64):

initial conditions are set along the same point boundary values describe an interval or range

OpenStudy (amistre64):

also, that looks like a burnolli equation with that y^4

OpenStudy (amistre64):

it most likely has some constraints for y=0, since you have to divide by y^4 to get ahead in it

OpenStudy (amistre64):

\[6y^{-4}y'-2y^{-3}=x\] z=y^-3 ; z^-1/3 = y y^-4 = z^-4/3 dz/dy = -3 y^-4 dz = -3 y^-4 dy \[6z^{-4/3}(\frac{z'}{-3z^{-4/3}})-2z=x\] \[-2z'-2z=x\] \[z'+z=-\frac{1}{2}x\]

OpenStudy (amistre64):

\[e^xz'+e^xz=\frac{1}{2}xe^x\] \[e^xz=\int \frac{1}{2}xe^x\ dx\] \[e^xz=\frac{1}{2}e^x(x-1)\] \[z=\frac{x-1}{2}=y^{-3}\]

OpenStudy (amistre64):

\[y=\sqrt[3]{\frac{2}{x-1}}\] x cannot be 1

OpenStudy (amistre64):

when x=0 we are to the left of 1; so our interval of validity is going to be at most: (-inf,1)

OpenStudy (amistre64):

which im pretty sure i just typed out for the practice since I dont see the asker anywhere lol

OpenStudy (amistre64):

since we are not solving for a specific solution other than when x=0; i think i did the math correctly :)

OpenStudy (amistre64):

well, i see where i dropped a negative tho; but that doesnt seem to change the outcome too much

OpenStudy (shayaan_mustafa):

@amistre64 can you tell me how did you recognize that it is Bernoulli equation?

OpenStudy (amistre64):

the power of y was a give away; normally we want to avoid any y parts and its derivatives raised to a power

OpenStudy (shayaan_mustafa):

OK. So are there any specific steps to remember in order to solve Bernoulli equations?

OpenStudy (amistre64):

get the y parts to the same side so its set up like a ordinary linear differential equation; then substitute your leftover ys with an appropriate z

OpenStudy (amistre64):

once everything is in terms of z; it is your normal ODE

OpenStudy (amistre64):

solve for z thru the regular processes; and then zince z = y^... evaluate the z for the y

OpenStudy (amistre64):

to make sure i didnt throw in an error roo bad :) \[e^xz'+e^xz=-\frac{1}{2}xe^x\] \[e^xz=-\int \frac{1}{2}xe^xdx\] \[e^xz=-\frac{1}{2}e^x(x-1)+C\] \[z=-\frac{1}{2}(x-1)+Ce^{-x}=y^{-3}\] \[-\frac{(x-1)+Ce^{-x}}{2}=y^{-3}\] \[-\frac{2}{(x-1)+Ce^{-x}}=y^{3}\] \[-\sqrt[3]{\frac{2}{(x-1)+Ce^{-x}}}=y\] when x=0; \[-\sqrt[3]{\frac{2}{(-1)+C}}=-2\] \[\frac{2}{(-1)+C}=8\] \[16=-1+C;\ C=17\] \[-\sqrt[3]{\frac{2}{(x-1)+17e^{-x}}}=y\] this of course goes bad at" \[x-1+17e^{-x} = 0\] which of course can be figured out ...later maybe

OpenStudy (amistre64):

i am most likely outa practice since the wolf come up with a slightly altered solution http://www.wolframalpha.com/input/?i=6y%27-2y%3Dxy%5E4%2C+y%280%29%3D-2